OpenStudy (anonymous):

Solve by simulating the problem.   You have a 5-question multiple-choice test. Each question has four choices. You don’t know any of the answers. What is the experimental probability that you will guess exactly three out of five questions correctly? Type your answer below using complete sentences.

OpenStudy (anonymous):

@nader1

OpenStudy (anonymous):

I'll be right back! AFK

OpenStudy (anonymous):

you want answer or i wait you?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

@nader1 ok im back

OpenStudy (anonymous):

to get the experimental probability, you have to run the experiment.

OpenStudy (anonymous):

ok how do i do that?

OpenStudy (anonymous):

this is a binomial situation

OpenStudy (anonymous):

Consider the probability of success for each trial, the number of trials, the number of successes sought, and the number of combinations of getting that many successes.

OpenStudy (anonymous):

1 in 4?

OpenStudy (anonymous):

exactly 3 right out of 5 means 3 right, 2 wrong and (5/2) ways to do it

OpenStudy (anonymous):

Number of ways to get 3 out of 5 = 5C3 = 5C2 = 10, so the 10 in right.

OpenStudy (anonymous):

so its 5/2 ways to do it

OpenStudy (anonymous):

... *is right. 3 successes = (.25)^3 with a 25% chance of getting any question correct at random. 2 failures = (.75)^2 for similar reasons as above.

OpenStudy (anonymous):

See http://stattrek.com/probability-distributions/binomial.aspx?Tutorial=Stat for derivation of the binomial formula.

OpenStudy (anonymous):

I think you have everything you need; just multiply it all together.

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

There are ten ways to get 3 out of 5. (use combinations formula 5choose3) To get 3 successes with a 25% chance each, that's 0.25 × 0.25 × 0.25 = (0.25)^3 To also get the 2 failures with a 75% chance: 0.75 × 0.75 = (0.75)^2 Altogether, the probability is 10 × (0.25)^3 ×(0.75)^2.

OpenStudy (anonymous):

so thats the probability...thanks

OpenStudy (anonymous):

this one was reallly confusing..:)

OpenStudy (anonymous):

Makes sense to me. 10 ways to have 3 successes and 2 failures multiplied by the probability of having 3 successes multiplied by the probability of having 2 failures. See also http://stattrek.com/probability/probability-rules.aspx?Tutorial=Stat

OpenStudy (anonymous):

Follow those links to that website I posted and study until you understand why the probability is computed that way

OpenStudy (anonymous):

good teaching :P?

OpenStudy (anonymous):

ok thanks nader brb

OpenStudy (anonymous):

you welcome :)

OpenStudy (anonymous):

dont forget medal ^^

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

anytime :) i hope i help you good